# Inhomogeneous Abelian Chern-Simons Higgs Model with New Inhomogeneous BPS Vacuum and Solitons

### v1

@@ Abstract
We study an inhomogeneous U(1) Chern-Simons Higgs model with a magnetic impurity in the BPS limit. The potential is sextic with both broken and unbroken phases, but its minimum varies spatially depending on the strength of the impurity. While the system lacks translation symmetry, it admits a supersymmetric extension. Depending on the sign of the impurity term, it has either a BPS sector or an anti-BPS sector (but not both), which satisfies the Bogomolny equations. The vacuum configuration of the broken phase is not simply determined by the the minimum of the potential since it is no longer constant, but it becomes a nontrivial function satisfying the Bogomolny equations. Thus, the energy and angular momentum densities of the vacuum locally have nonzero distributions, although the total energy and angular momentum remain zero. As in the homogeneous case, the theory supports various BPS soliton solutions, including topological and nontopological vortices and Q-balls. The vorticities as well as the U(1) charges are exclusively positive or negative. For a Gaussian type impurity as a specific example, we obtain rotationally symmetric numerical solutions and analyze their detailed properties.
@@ 1 Introduction
The rest of the paper is organized as follows. In section 2, we introduce the ICSH model with a magnetic impurity term and derive Bogomolny equations. We also identify the supersymmetric Lagrangian and construct the supersymmetry algebra. In section 3, we study the vacuum configurations of ICSH model in detail and obtain explicit solutions numerically for Gaussian impurities. In section 4, we obtain various topological as well as nontopological soliton solutions. We conclude in section 5.
@@ 2 BPS Limit of Inhomogeneous Chern-Simons Higgs Model
In $(2+1)$ dimensions, abelian Chern-Simons Higgs model is described by the Lagrangian density
@@
where the last term comes from $\Delta\mathscr{L}$ in (2.7). This theory has been first introduced in [28].
@@
Though the Poincare symmetry is explicitly broken, BPS nature of the theory suggests that it still has a supersymmetric extension. In fact, it is precisely given by the $\mathcal{N}=2$ supersymmetric homogeneous abelian Chern-Simons Higgs model [26] modified by $\Delta\mathscr{L}$ in (2.7) without further correction,
@@ 3 Inhomogeneous BPS Vacuum
One can obtain a similar behavior for the angular momentum $J$ in (2.23). Namely, the angular momentum density $\mathcal{J}=\epsilon^{ij}x_{i}T^{0j}$ is locally nonzero but its integral over the space is zero, in accordance with what we have shown above. This can be seen from the expression (2.26). Since $|\phi|$ is an increasing function of $r$ for the vacuum solution, the sign of $\mathcal{J}$ is completely determined by the factor $|\phi|^{2}-v^{2}(\boldsymbol{x})$. Thus $\mathcal{J}$ should be positive near the origin but becomes negative for large $r$. In Figure 2, we plot $\mathcal{J}$ for $\alpha^{-2}=20$ and $\beta=1$. At the center $\mathcal{J}$ vanishes. Then, it is surrounded by a ring with positive $\mathcal{J}$, which is, in turn, compensated by negative $\mathcal{J}$ in the outer region, as it should be. Therefore, the vacuum consists of regions rotating in different directions. Comparing with the energy density $T_{00}$ in Figure 1, we see that regions with positive/negative $T_{00}$ do not coinside with those with positive/negative $\mathcal{J}$. Considering that the inhomogeneous source $\sigma(\boldsymbol{x})$ does not directly contribute to $T_{0i}$, it is rather intriguing how the fields conspire to keep the vacuum spinless by locally adjusting the angular momentum density, while maintaining zero energy at the same time. It should be due to a nontrivial role of the Chern-Simons gauge field. In fact this kind of phenomenon does not occur in the inhomogeneous abelian Higgs model [32].
@@
One can obtain a similar result when the size parameter $\alpha$ is changed. See Figure 4.
Figure 4: (a) Amplitudes of the scalar field of the inhomogeneous broken vacuum for a Gaussian impurity (3.37) with $\beta=1$ and various $\alpha$. (b) Energy densities $T_{00}$ of the vacuum in unit of $\kappa m^{3}$.
@@ 5 Conclusions
In this paper, we have studied the Chern-Simons Higgs model with a magnetic impurity which allows supersymmetric extension with a suitable choice of the impurity term. Compared to the homogeneous theory without an impurity, the number of supersymmetries is reduced from $\mathcal{N}=2$ to $\mathcal{N}=1$. Since the translation symmetry is broken, vacuum solutions need not be constant in this theory. We showed that, in addition to the trivial vacuum $\phi=0$ in the symmetric phase, the vacuum in the broken phase has a nontrivial profile satisfying the Bogomolny equations. It has nontrivial energy density, magnetic field and angular momentum density while its energy, magnetic flux and angular momentum remain all zero. Depending on the sign of the impurity term added to the Lagrangian, the Bogomolny equations only have vortex solutions with either positive vorticities or negative vorticities, but not both.
@@
In the homogeneous Chern-Simons Higgs model, there are also domain wall solutions [27] which connect the symmetric vacuum and the broken vacuum. We expect that such solutions also exist in inhomogeneous theories as long as the vacuum structure does not change. In this paper, we investigated theories with a U(1) Chern-Simons gauge field. One can consider, instead, theories with a Maxwell term [15, 16, 32] or both. The gauge field can also be nonabelian [12]. We will report the results on these issues in separate publications.
@@ Acknowledgement
We would like to thank Seungjun Jeon for useful discussions. This work was supported by the National Research Foundation of Korea(NRF) grant with grant number NRF-2022R1F1A1074051 (C.K.), NRF-2022R1F1A1073053 (Y.K.), and RS-2023-00249608, NRF-2019R1A6A1A10073079 (O.K.).

### v2

@@ Abstract
We study an inhomogeneous U(1) Chern-Simons Higgs model with a magnetic impurity in the BPS limit. The potential is sextic with both broken and unbroken phases, but its minimum varies spatially depending on the strength of the impurity. While the system lacks translation symmetry, it admits a supersymmetric extension. Depending on the sign of the impurity term, it has either a BPS sector or an anti-BPS sector (but not both), which satisfies the Bogomolny equations. The vacuum configuration of the broken phase is not simply determined by the the minimum of the potential since it is no longer constant, but it becomes a nontrivial function satisfying the Bogomolny equations. Thus, the energy and angular momentum densities of the vacuum locally have nonzero distributions, although the total energy and angular momentum remain zero. As in the homogeneous case, the theory supports various BPS soliton solutions, including topological and nontopological vortices and Q-balls. The vorticities as well as the U(1) charges are exclusively positive or negative. For a Gaussian type impurity as a specific example, we obtain rotationally symmetric numerical solutions and analyze their detailed properties. We also discuss the case of a delta-function impurity as the infinitely thin limit of the Gaussian impurity which shows some nontrivial feature of BPS Chern-Simons Higgs theory.
@@ 1 Introduction
We also consider the case of delta-function impurity as the infinitely thin limit of the Gaussian impurity. Due to the characteristic nature of the Chern-Simons field, the broken vacuum solution in this limit turns out to be ill-defined at the impurity point. Away from the point, however, the delta-function impurity has no effect on the BPS solutions and the solutions are exactly the same as those in homogeneous theory with no impurity.
The rest of the paper is organized as follows. In section 2, we introduce the ICSH model with a magnetic impurity term and derive Bogomolny equations. We also identify the supersymmetric Lagrangian and construct the supersymmetry algebra. In section 3, we study the vacuum configurations of ICSH model in detail and obtain explicit solutions numerically for Gaussian impurities and its limit to delta-function impurities. In section 4, we obtain various topological as well as nontopological soliton solutions. We conclude in section 5. There are two appendices. Appendix A discusses an off-shell formulation of the supersymmetric ICSH model where electric impurities are also briefly considered. In appendix B, we study another type of regularized delta-function impurity and show that the value of the scalar field at the impurity point depends on the regularization.
Note added: After submitting the paper, we became aware of the recent paper [29] which has a partial overlap with our work. It considers the same ICSH model with a Gaussian-type magnetic impurity and discusses some basic features of the vacuum in the broken phase as well as topological soliton solutions.
@@ 2 BPS Limit of Inhomogeneous Chern-Simons Higgs Model
In $(2+1)$ dimensions, abelian Chern-Simons Higgs (CSH) model is described by the Lagrangian density
@@
where the last term comes from $\Delta\mathscr{L}$ in (2.7). This theory has been first introduced in [28] and further studied in [29].
@@
Though the Poincare symmetry is explicitly broken, BPS nature of the theory suggests that it still has a supersymmetric extension. In fact, it is precisely given by the $\mathcal{N}=2$ supersymmetric homogeneous abelian CSH model [26] modified by $\Delta\mathscr{L}$ in (2.7) without further correction,
@@
Although we will mainly consider the system with magnetic impurities in this paper, it is also possible to add electric impurities or both to the system while keeping the BPS nature. We briefly discuss the case by considering the off-shell formulation of its supersymmetric inhomogeneous model in the appendix A.
@@ 3 Inhomogeneous BPS Vacuum
One can obtain a similar behavior for the angular momentum $J$ in (2.23). Namely, the angular momentum density $\mathcal{J}=\epsilon^{ij}x_{i}T^{0j}$ is locally nonzero but its integral over the space is zero, in accordance with what we have shown above. This can be seen from the expression (2.26). Since $|\phi|$ is an increasing function of $r$ for the vacuum solution, the sign of $\mathcal{J}$ is completely determined by the factor $|\phi|^{2}-v^{2}(\boldsymbol{x})$. Thus $\mathcal{J}$ should be positive near the origin but becomes negative for large $r$. In Figure 2, we plot $\mathcal{J}$ for $\alpha^{-2}=20$ and $\beta=1$. At the center $\mathcal{J}$ vanishes. Then, it is surrounded by a ring with positive $\mathcal{J}$, which is, in turn, compensated by negative $\mathcal{J}$ in the outer region, as it should be. Therefore, the vacuum consists of regions rotating in different directions. Comparing with the energy density $T_{00}$ in Figure 1, we see that regions with positive/negative $T_{00}$ do not coincide with those with positive/negative $\mathcal{J}$. Considering that the inhomogeneous source $\sigma(\boldsymbol{x})$ does not directly contribute to $T_{0i}$, it is rather intriguing how the fields conspire to keep the vacuum spinless by locally adjusting the angular momentum density, while maintaining zero energy at the same time. It should be due to a nontrivial role of the Chern-Simons gauge field. In fact this kind of phenomenon does not occur in the inhomogeneous abelian Higgs model [33].
@@
Now we consider the $\delta$-function limit of the impurity (3.3). With $\beta=4\alpha\eta$, the inhomogeneous function $\sigma$ in (3.3) is rewritten as
$\sigma(r)$ $=-4\eta\alpha v_{0}^{2}e^{-\alpha m^{2}r^{2}}$
$\equiv-\frac{\pi\eta\kappa^{2}}{v_{0}^{2}}\delta_{\alpha}(r),$ (3.10)
where $\delta_{\alpha}(r)$ becomes the two-dimensional delta function $\delta(\boldsymbol{x})$ in $\alpha\rightarrow\infty$ limit for fixed $\eta$. The vacuum equation (3.1) then leads to
$\nabla^{2}\ln|\phi|^{2}=\frac{4}{\kappa^{2}}|\phi|^{2}(|\phi|^{2}-v_{0}^{2})+\frac{4\pi\eta}{v_{0}^{2}}|\phi(\boldsymbol{x})|^{2}\delta_{\alpha}(r).$ (3.11)
Note that $|\phi(\boldsymbol{x})|^{2}$ is multiplied in front of the impurity function $\delta_{\alpha}(r)$, which is a characteristic feature of the BPS CSH model having the sextic potential (2.2). Without this factor, the delta-function limit of the last term in (3.11) is similar to the delta-function term in (2.34), which comes from the vortex points where the scalar field $\phi$ vanishes. Then, near the impurity point $r=0$, $\phi$ would tend to zero with some power of $r$ which depends on the coefficient of the delta function [^4] This is exactly the case in the inhomogeneous abelian Higgs model where the dynamics is governed by the Maxwell term [33]. In this model, the vacuum equation corresponding to (3.11) is $\nabla^{2}\ln|\phi|^{2}=2g^{2}(|\phi|^{2}-v_{0}^{2})+4\pi\eta\delta_{\alpha}(\boldsymbol{x}),$ (3.12) where $g$ is the coupling constant of the model. Thus, $\phi\sim r^{\eta}$ near the origin..(See section 4.) However, due to the factor $|\phi|^{2}$ in front of the delta function, this cannot happen because the vanishing $|\phi(0)|$ would kill the delta function itself. So one should not naively assume that the delta-function impurity would induce a zero of $\phi$.
To see the effect of the delta-function impurity, we obtain numerical solutions of the vacuum by changing the size parameter $\alpha$ and the depth parameter $\beta=4\alpha\eta$ for fixed $\eta$. Figure 4 shows that, as we increase $\alpha$, the dip of $|\phi|$ at the origin becomes narrower, as expected, but also more shallow at the same time. It seems clear from the figure that, as $\alpha\rightarrow\infty$, the vacuum solution simply becomes the constant configuration $|\phi|=v_{0}$ throughout the space including the origin. Although the scalar field $\phi$ does not directly show the effect of the delta function impurity, it appears in the the energy density as shown in Figure 4 where delta function term can be identified. This can also be seen from the Bogomolny equation (2.18). With $|\phi|\simeq v_{0}$, the magnetic field behaves as
$B\simeq\frac{2v_{0}^{2}}{\kappa^{2}}\sigma(\boldsymbol{x})=2\pi\eta\delta_{\alpha}(r).$ (3.13)
Nevertheless, the total magnetic flux as well as the total energy should remain zero since this is the vacuum solution. The localized flux and energy at the origin are compensated by the slight excess of the flux and energy in the rest of the space, which eventually go to zero in the delta-function limit.
Figure 4: (a) Amplitudes of the scalar field of the inhomogeneous broken vacuum for different $\alpha$ and $\beta=4\alpha\eta$ with $\eta=1$ fixed. (b) Energy densities $T_{00}$ of the vacuum in unit of $\kappa m^{3}$.
There is a subtle point in this result. One may argue that the constant solution $|\phi|=v_{0}$ seem to contradict the existence of the delta function in (3.11), since it does not vanish with nonzero $|\phi(0)|$ and should have some effect on the solution. What we really find here is that the delta-function limit of the solution is not well-defined at the origin; it becomes singular and the value of $\phi(0)$ and/or $\phi^{\prime}(0)$ depends on the limiting procedure. With the Gaussian regularization of the delta function adopted here, $\phi^{\prime}(0)$ becomes ill-defined, although $\phi(0)$ itself goes to $v_{0}$. On the other hand, it can be shown that a different result is obtained if a different regularization is used. We verify this explicitly in Appendix B. The difference, however, exists just at the single point, which should not change any physical quantities.
@@ 4.1 Topological vortices
We conclude this subsection with a brief comment on the delta function impurity considered at the end of section 3. Since vortex solutions behave as $|\phi|\sim r^{n}$ near the origin, the presence of the impurity term $|\phi|^{2}\delta(\boldsymbol{x})$ should not alter the solutions of the homogeneous theory as long as the nonzero vorticity $n$ is taken into account. This is true for both topological and nontopological vortices discussed in subsection 4.3. The only possible exception is the Q-ball solutions in the next subsection, but this case is exactly the same as the vacuum case.
@@ 5 Conclusions
In this paper, we have studied the CSH model with a magnetic impurity which allows supersymmetric extension with a suitable choice of the impurity term. Compared to the homogeneous theory without an impurity, the number of supersymmetries is reduced from $\mathcal{N}=2$ to $\mathcal{N}=1$. Since the translation symmetry is broken, vacuum solutions need not be constant in this theory. We showed that, in addition to the trivial vacuum $\phi=0$ in the symmetric phase, the vacuum in the broken phase has a nontrivial profile satisfying the Bogomolny equations. It has nontrivial energy density, magnetic field and angular momentum density while its energy, magnetic flux and angular momentum remain all zero. Depending on the sign of the impurity term added to the Lagrangian, the Bogomolny equations only have vortex solutions with either positive vorticities or negative vorticities, but not both.
@@
We also considered a delta-function impurity as a thin limit of the Gaussian impurity. Thanks to the special feature of the BPS CSH model, it does not alter the solutions except at the impurity point as far as the BPS solutions are concerned. It would be interesting to investigate further any implications of the result.
In this paper, we considered only magnetic impurities but we can also include electric impurities as well as both electric and magnetic impurities at the same time without breaking the BPS nature as discussed in the Appendix A. It would then be a natural generalization to investigate inhomogeneous models with these impurities. In this regards, we note that CSH with constant background charge density was studied some time ago [34] which may be considered as a constant electric impurity in this setup.
In the homogeneous CSH model, there are also domain wall solutions [27] which connect the symmetric vacuum and the broken vacuum. We expect that such solutions also exist in inhomogeneous theories as long as the vacuum structure does not change. In this paper, we investigated theories with a U(1) Chern-Simons gauge field. One can consider, instead, theories with a Maxwell term [15, 16, 33] or both. The gauge field can also be nonabelian [12]. We will report the results on these issues in separate publications.
@@ Acknowledgement
We would like to thank Seungjun Jeon for useful discussions and J. G. F. Campos for informing us of the paper [29]. This work was supported by the National Research Foundation of Korea(NRF) grant with grant number NRF-2022R1F1A1074051 (C.K.), NRF- 2022R1F1A1073053 (Y.K.), RS-2019-NR040081 (Y.K., O.K.), and RS-2023-00249608 (O.K.).
Appendix A Off-Shell Formulation
We discuss the inhomogeneous supersymmetric CSH model in the off-shell formulation [35, 36]. The homogeneous $\mathcal{N}=2$ supersymmetric Lagrangian is given by
$\mathscr{L}_{\textrm{off-shell}}=$ $\,\frac{\kappa}{2}\epsilon^{\mu\nu\rho}A_{\mu}\partial_{\nu}A_{\rho}+i\frac{\kappa}{2}\bar{\chi}\chi-\kappa ND-\overline{D_{\mu}\phi}D^{\mu}\phi+i\bar{\psi}\gamma^{\mu}D_{\mu}\psi$
$+(|\phi|^{2}-v_{0}^{2})D-N^{2}|\phi|^{2}+|F|^{2}-i(\bar{\psi}\chi\phi+\bar{\chi}\psi\bar{\phi})-iN\bar{\psi}\psi.$ (A.1)
It can easily be shown that it reduces to the on-shell Lagrangian (2.27) after elimination of the auxiliary fields.
Motivated by the fact that an abelian gauge field gives rise to a global U(1) current via $\tilde{j}_{\mu}=\frac{1}{2}\epsilon_{\mu\nu\rho}F^{\nu\rho},$ we may consider turning on an external vector field $\hat{A}_{\mu}$ which minimally couples to the current $\tilde{j}_{\mu}\hat{A}^{\mu}$. It then results in a BF interaction $\frac{1}{2}\epsilon^{\mu\nu\rho}\hat{A}_{\mu}F_{\nu\rho}$. After supersymmetrization, we obtain [15]
$\mathscr{L}_{\textrm{BF}}$ $=\frac{1}{2}\epsilon^{\mu\nu\rho}\hat{A}_{\mu}F_{\nu\rho}+\hat{D}N+\hat{N}D+\textrm{fermions}$
$=\hat{A}_{0}F_{12}+\hat{N}D+\hat{F}_{12}A_{0}+\hat{D}N+\textrm{fermions},$ (A.2)
where $\hat{N}$ and $\hat{D}$ denote the neutral scalar and the auxiliary field of the background vector multiplet, respectively. This term is invariant under the following supersymmetric variation, which is the off-shell version of (2.28),
$\delta A_{\mu}$ $=\frac{1}{2}(\bar{\eta}\gamma_{\mu}\chi+\bar{\chi}\gamma_{\mu}\eta),$
$\delta N$ $=\frac{1}{2}(\bar{\eta}\chi+\bar{\chi}\eta),$
$\delta D$ $=-\frac{i}{2}(\bar{\eta}\gamma^{\mu}\partial_{\mu}\chi-\partial_{\mu}\bar{\chi}\gamma^{\mu}\eta),$ (A.3)
provided that the parameter $\eta$ satisfies the condition $\gamma^{1}\eta=-is\gamma^{2}\eta$, where $s=\pm 1$, and
$\hat{A}_{0}$ $=-s\hat{N}\equiv s\sigma,$
$\hat{F}_{12}$ $=-s\hat{D}\equiv s\sigma_{e}.$ (A.4)
Then, (A.2) becomes
$\mathscr{L}_{\textrm{BF}}=\sigma(sB-D)+\sigma_{e}(sA_{0}-N)+\textrm{fermions}.$ (A.5)
In this Lagrangian, $\sigma D$ term plays the role of an inhomogeneous version of the Fayet–Iliopoulos term and nonzero $\sigma$ precisely corresponds to the magnetic impurity considered in the main part of the paper. On the other hand, nonzero $\sigma_{e}$ gives an electric impurity. In particular, if the impurity is point-like, e.g. $\sigma_{e}=\delta(\boldsymbol{x})$, it simply reduces to a familiar $1/2$ BPS Wilson line. Thus, the electric impurity term in (A.5) is possibly considered as a superposition of those Wilson lines. Note that both the electric and the magnetic impurities can be turned on simultaneously, while maintaining the supersymmetry.
Appendix B Notes on Delta-Function Impurity
In this appendix, we choose the following form of the regularized delta function at the origin as the magnetic impurity,
$\tilde{\delta}(r)=\frac{1}{\pi\epsilon^{2}}\theta(\epsilon-r),$ (B.1)
where $\epsilon$ is assumed to be very small. For $r<\epsilon$, the vacuum equation (3.11) becomes
$\nabla^{2}\ln|\phi|^{2}=\frac{4}{\kappa^{2}}|\phi|^{2}(|\phi|^{2}-v_{0}^{2})+\frac{4\eta}{\epsilon^{2}v_{0}^{2}}|\phi|^{2}.$ (B.2)
When $\epsilon$ is small, it is clear that the second term on the right hand side dominates and (B.2) reduces to the Liouville equation. It has a rotationally symmetric solution which is regular and nonvanishing at the origin,
$|\phi|^{2}=\frac{2c^{2}\epsilon^{2}v_{0}^{2}}{\eta(c^{2}-r^{2})^{2}},$ (B.3)
where $c$ is a constant. For $r>\epsilon$, the impurity term vanishes and we expect $|\phi|\simeq v_{0}$. In this region, the solution is given by the asymptotic behavior (3.8). Connecting the two solutions (B.3) and (3.8) by requiring the continuities of $|\phi|$ and $d|\phi|/dr$ at $r=\epsilon$, we can obtain the constants $b$ and $c$ in the $\epsilon\rightarrow 0$ limit,
$b\simeq\frac{1}{2\ln\frac{1}{m\epsilon}},\qquad c\simeq\left(\frac{8}{\eta}\epsilon^{2}\ln\frac{1}{m\epsilon}\right)^{1/4}.$ (B.4)
Then, in the $\epsilon\rightarrow 0$ limit, we have
$|\phi(0)|^{2}$ $=\frac{2\epsilon^{2}v_{0}^{2}}{\eta c^{2}}\rightarrow 0,$
$|\phi(r)|^{2}$ $\rightarrow v_{0}^{2}\qquad(r\neq 0).$ (B.5)
Therefore, in the presence of the magnetic delta-function impurity which is defined as the limit of the regularized function given by (B.1), the vacuum solution is the constant solution $|\phi|=v_{0}$ except at the impurity point where $\phi$ vanishes, which is different from the result of section 3 where $|\phi|=v_{0}$ including at the impurity point. This verifies that the value of the scalar field depends on the regularization, although physical quantities should not be affected by such a difference at the single point.